A diver exhales a bubble with a volume of 1250 mL at a pressure of 2.4 atm and a temperature of 15 °C. How many gas moles are in this bubble?
Answer #1

Given:

P = 2.4 atm

V = 1250.0 mL

= (1250.0/1000) L

= 1.25 L

T = 15.0 oC

= (15.0+273) K

= 288 K

find number of moles using:

P * V = n*R*T

2.4 atm * 1.25 L = n * 0.08206 atm.L/mol.K * 288 K

n = 0.1269 mol

Answer: 0.127 mol